OpenStudy (anonymous):

please help write this equation in slope-intercept form of the line that passes through the given points. (-1,2) and (0,0)

OpenStudy (shamil98):

First find the slope of the equation.

OpenStudy (shamil98):

\[\large \frac{y_2 -y_1}{ x_2-x_1 }\] is your slope equation.

OpenStudy (shamil98):

\[\large \frac{ 0-2 }{ 0-(-1) } = -2\] From there input your x and y coordinates and slope into y = mx +b and solve for b. 0 = 0(-2) + b b = 0 your resulting equation is \[\large y = -2x\]